How much energy is needed to turn 2 kg of water, at 100°c? Specific latent heat of vaporization of water =2300 000 J\kg.


1
Expert's answer
2021-02-25T11:26:03-0500

By definition, the amount of energy required to vaporize "m = 2kg" of water at 100°C is:


"Q = \lambda m"

where "\lambda = 2.3\times 10^6J/kg" is the specific latent heat of vaporization of water. Thus, obtain:


"Q = 2.3\times 10^6J/kg\cdot 2kg = 4.6\times 10^6J"

Answer. "4.6\times 10^6J" or "4.6MJ".